### v1

# arXiv:2603.01305v1[cs.CV] 01 Mar 2026

Abstract
Large multimodal models (LMMs) exhibit strong task generalization capabilities, offering new opportunities for zero-shot visual anomaly segmentation (ZSAS). However, existing LMM-based segmentation approaches still face fundamental limitations: anomaly concepts are inherently abstract and context-dependent, lacking stable visual prototypes, and the weak alignment between high-level semantic embeddings and pixel-level spatial features hinders precise anomaly localization. To address these challenges, we present AG-VAS (Anchor-Guided Visual Anomaly Segmentation), a new framework that expands the LMM vocabulary with three learnable semantic anchor tokens—[SEG], [NOR], and [ANO], establishing a unified anchor-guided segmentation paradigm. Specifically, [SEG] serves as an absolute semantic anchor that translates abstract anomaly semantics into explicit, spatially grounded visual entities (e.g., holes or scratches), while [NOR] and [ANO] act as relative anchors that model the contextual contrast between normal and abnormal patterns across categories. To further enhance cross-modal alignment, we introduce a Semantic-Pixel Alignment Module (SPAM) that aligns language-level semantic embeddings with high-resolution visual features, along with an Anchor-Guided Mask Decoder (AGMD) that performs anchor-conditioned mask prediction for precise anomaly localization. In addition, we curate Anomaly-Instruct20K, a large-scale instruction dataset that organizes anomaly knowledge into structured descriptions of appearance, shape, and spatial attributes, facilitating effective learning and integration of the proposed semantic anchors. Extensive experiments on six industrial and medical benchmarks demonstrate that AG-VAS achieves consistent state-of-the-art performance in the zero-shot setting.
33footnotetext: Project Page: https://github.com/xiaozhen228/AG-VAS
1 Introduction
Figure 1: Comparison of zero-shot anomaly segmentation capabilities between LISA [21] and our AG-VAS.
Figure 2: AG-VAS is an effective and efficient LMM for zero-shot anomaly segmentation and understanding. (Left: raw image; Right: segmentation result.) The visualization results are shown across the following scenarios: (1)–(4) direct segmentation, (5)–(6) describe-then-segment, (7) segment-then-explain, and (8) segmentation within dialogue.
Zero-shot visual anomaly segmentation (ZSAS) aims to directly localize anomalous regions on the surfaces of unseen-category objects without the need for retraining. This capability is particularly important in data-scarce or privacy-sensitive scenarios, such as industrial defect localization [40] and medical image analysis [2].
Most existing ZSAS models [19, 53, 10, 33, 12, 34, 11] are built upon CLIP [35] due to its strong vision–language alignment capabilities. These models typically construct pairs of textual prompts representing normal and abnormal states, and align them with pixel-level patch embeddings to localize anomalies. Representative methods, such as AnomalyCLIP [53] and Bayes-PFL [33], further enhance the fine-grained anomaly perception via prompt learning strategies. However, constrained by the limited representation and understanding capabilities of CLIP, the ZSAS performance of these approaches appears to have reached a bottleneck. Moreover, they do not naturally produce binary segmentation masks, often requiring heuristic thresholding or empirical tuning, which hinders practical deployment. These limitations motivate exploring large multimodal models (LMMs) with stronger reasoning capacity and richer world knowledge [28, 13, 44, 54, 36].
Recent anomaly detection studies [17, 46, 18, 24] leveraging LMMs have demonstrated remarkable performance in high-level visual understanding and conversational interaction. However, most existing works (e.g. Anomaly-OV [46]) focus on enhancing image-level textual descriptions or anomaly classification accuracy, while precise pixel-level responses, such as binary segmentation masks, remain largely unexplored. In this work, we aim to develop an end-to-end ZSAS framework built upon the pretrained LMMs, capable of understanding complex segmentation instructions and directly outputting binary masks.
LISA [21] is the first general-purpose image segmentation work to employ an embedding-as-mask scheme, extending visual comprehension to pixel-wise segmentation by integrating LMMs with dedicated segmenters (e.g. SAM [20]). However, when applied to industrial or medical ZSAS tasks, it often fails to correctly identify anomalous regions, sometimes even confusing foreground and background. For instance, given the instruction "Please segment the anomalies in this image", LISA may produce masks that mislocalize the abnormal regions, as illustrated in Figure 1. We attribute these failures in ZSAS to two main factors: 1) anomalies or defects are highly abstract concepts without concrete semantic references—for example, unlike a well-defined object such as “apple”, anomalies can include “holes” or “scratches”, making them difficult to map from textual descriptions; 2) the segmenter’s pixel-level features are not well aligned with the LMM’s visual-textual embedding space, limiting effective anomaly localization.
To address the above issues, we propose AG-VAS, an end-to-end ZSAS framework equipped with a set of learnable anchor tokens that encode contextual and world knowledge about anomalies. These anchors serve as semantic bridges between the LMM and the segmenter, enabling effective instruction-driven anomaly segmentation. Specifically, we extend the LMM vocabulary with special anchor tokens: the absolute semantic anchor [SEG] provides explicit cues regarding anomaly appearance, shape, and location, while the relative semantic anchors [NOR] and [ANO] capture contextual contrasts between normal and abnormal regions across diverse categories. To enhance cross-modal correspondence, we further introduce a Semantic–Pixel Alignment Module (SPAM), which aligns high-level semantic embeddings from the LMM with fine-grained pixel features from the segmenter. Finally, an Anchor-Guided Mask Decoder (AGMD) leverages these aligned representations for anchor-consistent mask generation, achieving binarized segmentation results.
Figure 3: The pipeline of AG-VAS. We design two types of semantic anchors: the absolute anchor [SEG], which injects world and contextual knowledge about the appearance, structure, and location of anomalies (e.g., holes, scratches) into the segmentation process, and the relative normal/anomalous anchors [NOR]/[ANO], which guide the model to understand anomalies in relation to normal patterns. All parameters except those for the semantic/pixel image encoder, visual projector and LLM are trainable.
To further enrich the model’s understanding of anomaly-related semantics, we construct Anomaly-Instruct20K, an instruction-tuning dataset specifically designed for segmentation tasks. Unlike conventional instruction datasets focused on dialogue or visual question answering, Anomaly-Instruct20K injects structured world knowledge of anomalies into the LMM through diverse descriptions of defect appearance, shape, and spatial context. This enables the anchor tokens to learn precise and discriminative semantic representations. Moreover, its instruction-based design allows AG-VAS to develop interactive segmentation behaviors and flexibly respond to various user prompts. Figure 2 demonstrates that AG-VAS produces high-quality masks under different segmentation instructions. In summary, our contributions are:
  • We propose AG-VAS, an anchor-guided framework for ZSAS that directly produces binary masks through instruction-driven segmentation. It introduces absolute and relative semantic anchors to bridge the LMM and segmenter, and incorporates two modules: SPAM for aligning LMM and segmenter features, and AGMD for generating accurate, anchor-consistent masks.
  • We construct Anomaly-Instruct20K, a segmentation-oriented instruction-tuning dataset that injects structured world knowledge of anomalies into the LMM, enhancing its understanding of defect-related semantics.
  • We achieve state-of-the-art performance on multiple industrial and medical anomaly detection benchmarks, demonstrating the strong generalization ability and practical applicability of AG-VAS.
2 Related Work
2.1 Large Multimodal Models
Recent advances in LMMs can be broadly categorized into two paradigms: comprehension-oriented and perception-oriented approaches. Comprehension-oriented methods focus on enhancing visual understanding by aligning visual encoders with language models and improving representation quality through progressive training or multi-scale feature design [23, 1, 16, 15, 25, 28, 43]. Representative models such as BLIP-2 [23], Flamingo [1], LLaVA [28], and Qwen-VL [43] have substantially advanced cross-modal reasoning and instruction-following capabilities. More recently, techniques like chain-of-thought reasoning [29, 32] and mixture-of-experts architectures [47, 27] have further improved LMMs’ ability to handle complex visual comprehension tasks. In contrast, perception-oriented LMMs emphasize fine-grained visual grounding, aiming to adapt language models for pixel-level localization tasks [21, 36, 38, 26, 6, 45, 51]. A promising direction is the embedding-as-mask paradigm, first introduced by LISA [21] and later extended by PixelLM [36] and PaDT [38]. While these methods achieve strong performance in general object segmentation, they remain suboptimal when applied to the ZSAS task.
2.2 Zero-Shot Anomaly Segmentation
Recent studies on ZSAS have highlighted the potential of foundation models such as CLIP [35] for localizing anomalies in unseen categories. Most existing methods [19, 10, 53, 11, 34, 33, 31, 55] leverage CLIP’s strong vision–language alignment and improve fine-grained anomaly understanding through feature adaptation or prompt learning. Early works like APRILN-GAN [11] and CLIP-AD [12] introduce learnable linear layers to project image patch features into the joint embedding space. Later methods strengthen cross-category ZSAS performance through various prompt-learning designs, including static learnable prompts (AnomalyCLIP [53]), visual-context prompts (VCP-CLIP [34]), hybrid static–dynamic prompts (AdaCLIP [10]), and prompt distribution learning (Bayes-PFL [33]). Several LMM-based methods have recently been proposed for zero-shot tasks [46, 17, 24, 18], such as Anomaly-OV [46] and IAD-R1 [24]. However, they primarily focus on image-level anomaly understanding and classification rather than segmentation. In contrast to the above methods, this work aims to develop an instruction-following ZSAS model based on LMMs that directly outputs binarized segmentation masks, which is more practical in real industrial and medical scenarios.
3 Method
Following the recent ZSAD setting [53, 33], our model is trained under supervision on seen categories $C^{s}$ from an auxiliary dataset including our curated Anomaly-Instruct20K, and directly tested on unseen categories $C^{u}$ collected from industrial and medical domains. Importantly, $C^{s}$ and $C^{u}$ are disjoint, and the datasets used for training and evaluation exhibit substantial domain gaps.
3.1 Model Design
Framework overview. As illustrated in Figure 3, AG-VAS is an anchor-guided framework that integrates absolute and relative semantic anchors into an instruction-driven segmentation pipeline. It consists of four key components: 1) a frozen semantic image encoder $\mathcal{F}_{s}$ and pixel image encoder $\mathcal{F}_{p}$; 2) a large language model $\mathcal{F}_{\text{LLM}}$; 3) a Semantic–Pixel Alignment Module (SPAM); and 4) a lightweight Anchor-Guided Mask Decoder $\mathcal{D}$.
Given an input image $x_{\text{img}}$ and a text query $x_{\text{txt}}$, the frozen encoders first extract semantic features $f_{s}$ and pixel-level features $f_{p}$. These features are processed by the LLM together with the anchor tokens [NOR]/[ANO] and [SEG]. The mask decoder $\mathcal{D}$ then decodes the final-layer embedding corresponding to each anchor token to produce the binary segmentation mask $\mathbf{M}$. SPAM enhances this process by aligning the LLM’s semantic representations with the pixel-level features, enabling the anchor embeddings to effectively guide mask generation. In the following, we describe each component in detail.
Design of Semantic Anchors. In LMM-based ZSAS, we define semantic anchors as bridges that connects the high-level semantic representations produced by the LMM with the mask decoder. The key motivation is to enable the model to capture fine-grained anomaly-related semantics through specific anchor embeddings in the LMM output space. From the perspective of a human inspector, locating anomalies in unseen categories generally relies on two complementary reasoning processes: 1) leveraging prior world knowledge and contextual understanding of common defect types (e.g., holes, cracks, scratches) to recognize what anomalies might look like and where they typically appear; and 2) contrasting candidate regions against surrounding normal areas to identify inconsistencies or irregular patterns.
Inspired by these two reasoning modes, we introduce two types of semantic anchors: the absolute anchor [SEG], which encodes world knowledge and contextual cues regarding the appearance, structure, and location of potential anomalies, and the relative anchors [NOR]/[ANO], which guide the model to perceive anomalies in relation to normal patterns. These anchors are implemented as special tokens that are inserted into the text sequence and added in the LMM’s vocabulary space, allowing them to learn task-specific embeddings that differ from those obtained through conventional prompt tuning. During inference, AG-VAS mainly supports two segmentation modes that differ in whether the reasoning process is explicitly externalized:
(1) Implicit Mode. The model directly segments anomalies by leveraging internalized contextual and world knowledge. In this work, this corresponds to the Direct Segmentation scenario. For example:
User: Please segment the anomalies in this image.
Assistant: Sure, it is [NOR][ANO][SEG].
(2) Explicit Mode. The model externalizes its reasoning process either before or after mask prediction, such as describing anomalies before segmentation (Describe-then-Segment) or explaining the predicted mask (Segment-then-Explain). As an example, in the Describe-then-Segment process:
User: What anomalies are present in this image? Please describe the anomalies and then output the segmentation map. Assistant: The anomaly is a small hole located on the upper part of the hazelnut, as indicated by [NOR][ANO][SEG].
Both modes rely on identical anchor representations and world knowledge grounding; they differ only in whether the description or reasoning process is explicitly articulated. Unless otherwise specified, we adopt the Direct Segmentation mode for quantitative evaluation.
Semantic–Pixel Alignment Module. The introduction of semantic anchors bridges the high-level semantic understanding of LMMs to pixel-level segmentation. However, a gap exists between the high-level semantic embedding space of the LMM and the low-level visual features of the pixel encoder, hindering the anchors’ alignment with spatially coherent visual cues. Therefore, we propose the SPAM, which facilitates the alignment of pixel-level image features with the LMM’s embedding space through cross-modal interaction.
Figure 4: (a) An example from our Anomaly-Instruct20K dataset. (b) An example of instruction-tuning with dynamic fusion of templates during the actual training process. Note that only one specific template is shown here, with further details provided in the Appendix.
Specifically, given an image $x_{\text{img}}\in\mathbb{R}^{3\times h\times w}$, the semantic image encoder and the pixel image encoder are first used separately to extract visual features. The extracted features are then mapped to the same embedding dimension using a simple linear layer:
$f_{s}=Linear_{s}(\mathcal{F}_{s}(x_{\text{img}}))\quad f_{p}=Linear_{p}(\mathcal{F}_{p}(x_{\text{img}}))$ (1)
where $\mathcal{F}_{s}$ and $\mathcal{F}_{p}$ are semantic and pixel encoder, respectively. To facilitate better interaction between the anchors and the mask decoder, we employ a cross-modal attention module, where the semantic features extracted by the semantic image encoder serve as queries $\mathbf{Q}$ to attend to pixel embeddings ($\mathbf{K,V}$):
$f_{align}=MHCA(\mathbf{Q}=f_{s},\mathbf{K}=f_{p},\mathbf{V}=f_{p})$ (2)
where $MHCA$ denotes the multi-head cross attention [41]. Subsequently, the image embeddings $f_{s}$ from the semantic encoder, the aligned embeddings $f_{align}$, and the text embeddings $f_{\text{txt}}$ are concatenated and jointly fed into the large language model, which outputs a text response $\hat{y}_{\text{txt}}$:
$\hat{y}_{\text{txt}}=\mathcal{F}_{\text{LLM}}(Concat(f_{s},f_{align},f_{\text{txt}}))$ (3)
where $f_{\text{txt}}$ is a tokenized version of the input text query $x_{\text{txt}}$. The LLM output $\hat{y}_{\text{txt}}$ contains our designed anchor tokens [NOR],[ANO],[SEG]. We extract the LLM last-layer embeddings corresponding to these tokens and map them into the mask decoder space as $h_{nor},h_{ano},h_{seg}$ using a shared Token Refiner $\mathcal{F}_{tr}$, which consists of two linear layers.
Anchor-Guided Mask Decoder. Given the refined anchor embeddings $h_{\text{nor}},h_{\text{ano}},h_{\text{seg}}$ and the pixel embeddings $f_{p}$, the proposed AGMD aims to generate absolute and relative segmentation masks guided by semantic anchors. We first concatenate the anchor embeddings with three learnable query tokens $\mathbf{T}_{\text{learn}}=\{t_{\text{nor}},t_{\text{ano}},t_{\text{seg}}\}$ to form the decoder input:
$\mathbf{Z}_{0}=[t_{\text{nor}},t_{\text{ano}},t_{\text{seg}},h_{\text{nor}},h_{\text{ano}},h_{\text{seg}}].$ (4)
The decoder performs $L$ layers of bidirectional cross-attention between the initial latent feature $\mathbf{Z}_{0}$ and pixel embeddings $f_{p}$, resulting in the final output $\mathbf{Z}_{L}$ and updated pixel embeddings $f_{p}^{\prime}$. The process is described as:
$\mathbf{Z}_{L},f_{p}^{\prime}=\text{BiAttn}^{L}(\mathbf{Z}_{0},f_{p})$ (5)
where $\text{BiAttn}^{L}(\cdot)$ refers to applying a sequence of $L$ bidirectional cross-attention blocks, iteratively updating both $\mathbf{Z}_{0}$ and $f_{p}$, and following the design principles of SAM [20]. After the final layer, we select the updated learnable tokens $\mathbf{T}^{\prime}_{\text{learn}}=\{t^{\prime}_{\text{nor}},t^{\prime}_{\text{ano}},t^{\prime}_{\text{seg}}\}$ from $\mathbf{Z}_{L}$ to compute segmentation masks. More architectural details of AGMD are provided in the Appendix.
For the absolute semantic anchor [SEG], a foreground probability map is produced via a sigmoid activation $\sigma(\cdot)$:
$P_{\text{seg}}=\sigma(t^{\prime}_{\text{seg}}f_{p}^{\prime\top}),$ (6)
while for the relative anchors [NOR] and [ANO], a contrastive softmax function is applied to produce normal–abnormal probability maps:
$P_{\text{nor}},P_{\text{ano}}]=\text{Softmax}([t^{\prime}_{\text{nor}},t^{\prime}_{\text{ano}}]f_{p}^{\prime\top}).$ (7)
The final anomaly map is obtained by fusing the absolute and relative maps:
$P=\alpha\cdot P_{\text{seg}}+(1-\alpha)\cdot P_{\text{ano}},\quad\alpha=0.5$ (8)
Finally, the binary segmentation mask $\mathbf{M}$ is produced by thresholding the fused map at 0.5. This design allows the decoder to transform abstract anchor semantics into pixel-level representations under the guidance of both absolute ([SEG]) and relative ([NOR]/[ANO]) anchors. Detailed implementation and architecture specifications of AGMD are provided in the Appendix.
3.2 Training Objectives
The training of AG-VAS involves two complementary objectives:
1) Textual Autoregressive Loss. All tokens output by the LLM (including the anchor tokens) are supervised with the standard autoregressive cross-entropy loss:
$\mathcal{L}_{\text{txt}}=-\sum_{t}\log P(y_{t}\mid y_{<t},x_{\text{img}},x_{\text{txt}})$ (9)
where $y_{t}$ denotes the ground-truth token at step $t$, and $x_{\text{img}}$, $x_{\text{text}}$ are the input image and text query, respectively.
2) Segmentation Loss. For the predicted masks from the Anchor-Guided Mask Decoder, we unify the loss for all anchors as a combination of Binary Cross-Entropy (BCE) loss and Dice loss:
$\mathcal{L}_{\text{seg}}=\sum_{c}\Big(\lambda_{bce}\text{BCE}(P_{c},M_{c})+\lambda_{dic}\text{Dice}(P_{c},M_{c})\Big)$ (10)
where $c\in\{\text{SEG},\text{NOR},\text{ANO}\}$$P_{c}$ is the predicted mask for anchor $c$, and $M_{c}$ is the corresponding ground-truth mask. For the relative normal anchor, the ground-truth mask is the complement of the abnormal mask. The weights $\lambda_{bce}$ and $\lambda_{dic}$ are set to 0.5 and 2.0 empirically.
The total training objective is then defined as:
$\mathcal{L}=\mathcal{L}_{\text{txt}}+\mathcal{L}_{\text{seg}}$ (11)
Table 1: Comparison with existing state-of-the-art methods. The best results are marked in red, while the second-best methods are indicated in blue. The superscript * indicates that these methods were retrained on the anomaly segmentation dataset used in our AG-VAS.
Method Base Model Industrial Datasets Medical Datasets
MVTec-AD KSDD2 RSDD ISIC ColonDB ClinicDB
CLIP-Based (AP, F1-Max, $\text{IoU}_{\text{ano}}$) WinCLIP [19] CLIP(ViT-B-16-240) (18.0, 24.3, 0.0) (17.1, 24.6, 0.0) (2.1, 6.6, 0.1) (62.4, 64.1, 0.24) (14.3, 21.0, 0.0) (19.4, 27.3, 0.0)
APRIL-GAN [11] CLIP(ViT-L-14-336) (40.8, 43.3, 25.1) (61.6, 62.6, 41.8) (30.6, 40.2, 24.3) (69.8, 69.3, 12.8) (23.2, 32.5, 14.4) (38.8, 44.4, 25.3)
AnomalyCLIP [53] CLIP(ViT-L-14-336) (34.5, 39.1, 22.6) (41.8, 50.3, 27.6) (19.1, 30.6, 8.8) (74.4, 70.8, 51.0) (31.3, 39.8, 25.2) (42.2, 45.9, 30.0)
AnomalyCLIP* [53] CLIP(ViT-L-14-336) (35.2, 39.0, 22.4) (48.8, 53.6, 28.6) (25.5, 34.4, 8.5) (76.5, 71.4, 54.1) (34.2, 39.8, 22.7) (44.1, 47.3, 29.4)
Bayes-PFL [33] CLIP(ViT-L-14-336) (48.3, 49.1, 22.2) (73.7, 63.5, 25.4) (39.1, 45.2, 3.7) (84.6, 76.8, 60.0) (31.9, 39.2, 16.8) (53.2, 51.7, 21.0)
Bayes-PFL* [33] CLIP(ViT-L-14-336) (50.3, 50.4, 29.9) (74.9, 67.6, 39.5) (41.9, 46.9, 7.7) (81.9, 74.3, 52.1) (30.5, 38.1, 27.7) (47.6, 50.7, 34.4)
LMM-Based (AP, F1-Max, $\text{IoU}_{\text{ano}}$) PixelLM [36] LLaVA-13B (13.6, 19.6, 11.9) (42.6, 46.8, 32.4) (5.3, 9.4, 4.9) (69.8, 61.7, 48.9) (8.9, 16.8, 9.8) (13.0, 20.9, 14.6)
PaDT [38] Qwen2.5-VL-7B (6.8, 12.9, 16.4) (1.2, 2.9, 12.4) (0.9, 1.8, 1.7) (65.9, 68.3, 54.5) (35.4, 48.4, 39.1) (39.8, 51.4, 49.8)
LISA [21] LLaVA-7B (9.8, 16.9, 10.2) (19.5, 23.8, 18.2) (6.9, 12.5, 2.9) (68.1, 66.3, 54.6) (25.9, 33.1, 19.4) (50.0, 45.3, 36.4)
LISA [21] LLaVA-13B (14.6, 21.8, 13.7) (36.8, 39.8, 29.4) (5.2, 11.4, 2.6) (87.1, 78.1, 64.8) (35.3, 34.0, 26.2) (56.3, 52.9, 42.3)
LISA* [21] LLaVA-13B (37.6, 40.8, 31.8) (63.7, 63.8, 50.9) (52.1, 57.1, 35.2) (88.7, 82.4, 64.1) (40.9, 43.2, 29.5) (67.8, 62.7, 48.3)
LISA [21] LLaVA-OneVision-7B (32.4, 38.3, 32.2) (59.8, 65.8, 49.4) (31.6, 37.6, 22.6) (90.7, 81.6, 56.9) (29.1, 35.8, 14.4) (56.1, 54.4, 32.6)
LISA* [21] LLaVA-OneVision-7B (41.0, 44.1, 32.3) (69.4, 69.8, 48.4) (48.8, 54.9, 34.6) (84.9, 77.6, 50.1) (67.2, 62.9, 43.8) (81.0, 74.1, 59.8)
AG-VAS LLaVA-13B (40.5, 42.5, 37.0) (66.5, 65.4, 53.7) (54.3, 57.0, 35.9) (88.8, 82.7, 64.6) (43.7, 45.1, 32.2) (68.3, 65.5, 50.1)
AG-VAS LLaVA-OneVision-7B (51.0, 52.7, 44.8) (76.6, 74.7, 53.8) (56.2, 57.8, 36.9) (92.7, 84.9, 65.0) (70.7, 66.2, 58.2) (86.6, 79.2, 69.5)
3.3 Anomaly-Instruct20K Dataset
To improve the instruction-based segmentation ability of large multimodal models, we introduce a new dataset, Anomaly-Instruct20K, designed to inject world knowledge on normal and abnormal visual patterns into the LMM backbone. Different from prior datasets such as Anomaly-Instruct-125K [46], which mainly provide question–answer annotations for general anomaly understanding, our dataset bridges vision–language reasoning with pixel-level anomaly perception through instruction-driven segmentation supervision. Anomaly-Instruct20K is built using samples from three publicly available datasets: RealIAD [42], GoodsAD [50], and DTD-Synthetic [3].
We generate the dataset with the scientific multimodal model Inter-S1-241B [4]. For each anomalous sample, the model receives three visual inputs: 1) the raw image with bounding boxes, 2) the raw image overlaid with the ground-truth mask, and 3) a corresponding normal reference image. Conditioned on these inputs, Inter-S1 generates a structured annotation comprising five fields: Expectation, which describes the object’s ideal normal appearance; Observation, which details concrete visual deviations (e.g., location, shape, size, or texture irregularities); Diagnosis, which explains why these deviations break normal material or structural consistency; Summary, which converts the detected cues into a concise segmentation instruction and explicitly encodes spatial and appearance attributes necessary for grounding the anchor tokens [NOR][ANO][SEG]; and Explanation, which synthesizes all components into a coherent reasoning paragraph. An example of the generated annotations, along with the visual inputs, is shown in Figure 4(a).
During training, we generate final instructions by sampling templates from a library and combining them with dataset fields, covering four main instruction types: Direct Segmentation, Describe-then-Segment, Describe-then-Segment-Plus and Segment-then-Explain. An example of combining a template to generate a real instruction-tuning sample during the training process is shown in Figure 4(b). More detailed dataset construction procedures and system prompts are included in the Appendix.
4 Experiment
4.1 Experimental Setup
Datasets. The training of AG-VAS follows a multi-task joint learning strategy, leveraging datasets from two complementary domains: general object segmentation and anomaly segmentation. For general object segmentation, we adopt the setup from LISA [21], using datasets such as ADE20K [52] to provide comprehensive segmentation capabilities. For anomaly segmentation, we employ two sources of auxiliary training data: 1) the constructed Anomaly-Instruct20K, and 2) a random set of 20k images extracted from existing industrial anomaly segmentation datasets [30, 49, 37, 5, 3], referred to as Anomaly-Seg20K, used under direct segmentation instructions. To preserve the model’s dialogue and instruction-following capabilities, we include VQA datasets (LLaVA-150K [28] and WebAD [46]) during training. This joint training enables the absolute semantic anchor [SEG] to align with specific defect entities (e.g., holes, scratches) through contextual semantics and world knowledge, while the relative anchors [NOR] and [ANO] capture the semantic contrast between normal and anomalous regions. To evaluate the ZSAS performance of our model, we follow the protocol in [33] and conduct experiments on six real-world datasets, including three from the industrial domain (MVTec-AD [7], KSDD2 [9], RSDD [48]) and three from the medical domain (ISIC [14], ColonDB [39], ClinicDB [8]). More details about the datasets can be found in Appendix.
Evaluation Metrics. In this work, we focus on the quality of ZSAS predictions, particularly the binarized segmentation accuracy. We report pixel-level average precision (AP), the maximum F1 score at the optimal threshold (F1-Max), and the Intersection over Union of anomalous samples ($\text{IoU}_{\text{ano}}$). To further evaluate the model’s ability to reject normal images, we define $\text{IoU}_{\text{nor}}$: the model receives a score of 1 if it predicts an empty mask for a normal sample, and 0 otherwise. Both $\text{IoU}_{\text{ano}}$ and $\text{IoU}_{\text{nor}}$ are computed at the per-image level and then averaged over the corresponding subsets.
Implementation Details. By default, we adopt LLaVA-OneVision-7B [22] as our LMM backbone, equipped with a semantic image encoder (siglip-patch14-384), while SAM-ViT-H serves as the pixel image encoder for mask prediction. Training is implemented using DeepSpeed with the AdamW optimizer at a learning rate of 0.0003. During instruction tuning, the LLM is fine-tuned with LoRA at rank 16, using a global batch size of 80. The full training of the 7B model takes about 30 hours on four A100 GPUs with 40 GB memory each. More implementation details can be found in the Appendix.
Table 2: Comparison of the model’s ability to reject segmentation on normal samples on the MVTec-AD.
Method Base Model $\text{IoU}_{\text{ano}}$ $\text{IoU}_{\text{nor}}$ Average
PixelLM [36] LLaVA-13B 11.9 2.4 7.2
PaDT [38] Qwen2.5-VL-7B 16.4 0.0 8.2
LISA [21] LLaVA-7B 10.2 1.9 6.1
LISA [21] LLaVA-13B 13.7 6.3 10.0
LISA* [21] LLaVA-13B 31.8 43.5 37.7
LISA [21] LLaVA-OneVision-7B 32.2 4.0 18.1
LISA* [21] LLaVA-OneVision-7B 32.3 80.9 56.6
AG-VAS LLaVA-OneVision-7B 44.8 87.7 66.3
4.2 Comparison with State-of-the-art methods
Since LMM-based ZSAS is still in an early exploratory stage, directly comparable approaches are limited. Therefore, we include four representative CLIP-based ZSAS methods (WinCLIP [19], APRIL-GAN [11], AnomalyCLIP [53], and Bayes-PFL [33]) together with three segmentation methods that address referring segmentation tasks, specifically designed for natural scene understanding: PixelLM [36], PaDT [38], and LISA [21]. To ensure a fair comparison, we also retrained other methods (e.g., AnomalyCLIP, Bayes-PFL, and LISA) on the anomaly segmentation dataset used by our AG-VAS during training. More details regarding the comparison of these methods are shown in Appendix.
Quantitative Comparison. Table 1 presents a comparison of our AG-VAS with other state-of-the-art methods across various industrial and medical datasets. It can be observed that when using the default base model, LLaVA-OneVision, our ZSAS performance surpasses all other methods on every metric. CLIP-based methods, such as Bayes-PFL, achieve comparable AP and F1-Max scores relative to LMM-based approaches but perform poorly on metric $\text{IoU}_{\text{ano}}$, which measures true segmentation accuracy. This suggests that their anomaly localization results are ambiguous, leading to lower $\text{IoU}_{\text{ano}}$ after binarization. Furthermore, we observe that LLM-based methods demonstrate superior zero-shot generalization on medical datasets compared to CLIP-based approaches, particularly after fine-tuning on our carefully curated Anomaly-Instruct20K and Anomaly-Seg20K datasets. It is noteworthy that no medical images were included in the training dataset, demonstrating that LMMs can leverage pre-existing world knowledge for zero-shot generalization to unseen domains. Furthermore, the introduction of the anchor mechanism allows our method to more effectively align semantic and pixel-level features, leading to substantial improvements over other LMMs in leveraging contextual semantics during training.
Comparison of rejection performance. For LMM-based ZSAS methods, it is crucial to not only accurately segment anomalous regions but also to reject normal samples by predicting empty masks, even when the user provides instructions to segment anomalies. The experimental results in Table 2 show that most LMMs struggle in this regard, whereas our AG-VAS demonstrates strong rejection ability, achieving an $\text{IoU}_{\text{nor}}$ of 87.7% on normal samples. This indicates a substantial reduction in over-segmentation risk, while AG-VAS also maintains superior performance on anomalous regions, achieving an $\text{IoU}_{\text{ano}}$ of 44.8%.
4.3 Ablation
Table 3: Ablation on different components.
AP F1-Max $\text{IoU}_{\text{nor}}$ $\text{IoU}_{\text{ano}}$
Module Ablation w/o [SEG] 49.1 50.9 85.6 42.1
w/o [NOR][ANO] 47.2 49.7 52.1 39.5
w/o SPAM 46.5 48.0 70.6 41.4
Training Dataset Ablation w/o Anomaly-Instruct20K 48.4 50.0 85.2 39.9
w/o Anomaly-Seg20K 49.3 51.5 53.5 42.4
w/o General Segmentation 36.1 36.5 70.2 34.9
AG-VAS 51.0 52.7 87.7 44.8
In this subsection, ablation experiments on the MVTec-AD dataset are conducted to further investigate the impact of different settings on the proposed AG-VAS.
Influence of Different Components. Table 3 presents an ablation study evaluating the impact of various modules and training datasets on ZSAS performance. From the module ablations, it can be observed that removing the absolute semantic anchor [SEG] leads to a noticeable drop in all metrics, highlighting its crucial role in aligning anchor semantics with concrete defect entities. Excluding the relative anchors [NOR] and [ANO] significantly decreases $\text{IoU}_{\text{nor}}$, indicating their importance in capturing the semantic contrast between normal and anomalous regions. The removal of SPAM also causes performance degradation, suggesting its effectiveness in feature aggregation.
Regarding dataset ablations, training without Anomaly-Instruct20K leads to a noticeable decrease in performance, particularly on metrics reflecting anomaly segmentation. This indicates that instruction tuning with Anomaly-Instruct20K successfully injects domain-specific world knowledge into the model, enabling the anchors to better interpret and localize anomalies. Excluding Anomaly-Seg20K causes a larger drop in $\text{IoU}_{\text{nor}}$, showing that these auxiliary images are crucial for learning robust normal–abnormal contrasts. Omitting general segmentation data results in the most substantial decrease across all metrics, confirming that multi-domain joint training is essential for aligning absolute semantic anchors with visual features.
Influence of different segmentation modes during inference. Table 4 compares AG-VAS performance under different inference strategies. Describe-then-Segment refers to first generating a textual description of the target anomaly and then producing a segmentation mask. Describe-then-Segment-Plus prepends an additional sentence describing the normal appearance of the object to the input instruction before the anomaly description. This serves as an expert-provided prior, explicitly informing the model about the object’s typical state. Segment-then-Explain first predicts the mask and then generates an explanation, while Direct Segmentation predicts the mask directly from the instruction without intermediate text.
Compared with Direct Segmentation, Describe-then-Segment and Segment-then-Explain slightly reduce segmentation performance due to textual noise introduced during intermediate reasoning. However, these modes enable the model to better follow user instructions and support interactive reasoning, justifying a minor sacrifice in mask accuracy. Notably, Describe-then-Segment-Plus, by providing explicit knowledge of the object’s normal state, improves contextual understanding and partially mitigates the performance drop, leading to the best overall results among the descriptive modes.
5 Conclusion
This paper introduces AG-VAS, an anchor-guided framework for zero-shot visual anomaly segmentation that endows large multimodal models with learnable semantic anchors—absolute token [SEG] and relative tokens [NOR]/[ANO]—to jointly encode world knowledge about defect appearance, shape and location while simultaneously modeling the contextual opposition between normal and abnormal regions. By fusing high-level vision–language embeddings with fine-grained pixel features through the Semantic–Pixel Alignment Module, and decoding anchor-consistent masks via the bidirectional Anchor-Guided Mask Decoder, AG-VAS overcomes the lack of structured anomaly semantics and directly outputs accurate binary masks without requiring category-specific retraining. Extensive experiments on six challenging industrial and medical benchmarks demonstrate that AG-VAS establishes a new state-of-the-art, significantly outperforming both CLIP-based and recent LMM-based competitors. It shows strong rejection performance on normal images and demonstrates robust generalization across unseen domains, providing a practical solution for real-world zero-shot anomaly segmentation task.
Table 4: Ablation on different segmentation mode during inference.
Segmentation Mode AP F1-Max $\text{IoU}_{\text{nor}}$ $\text{IoU}_{\text{ano}}$
Describe-then-segment 49.3 51.3 58.8 43.6
Describe-then-segment-plus 51.9 53.0 90.2 45.1
Segment-then-explain 48.6 51.4 91.1 43.4
Direct Segmentation 51.0 52.7 87.7 44.8